# On The Collatz Conjecture: Topological and Ergodic Approach

### 03 Mar 2026 v3

@@ Abstract.
We study the Collatz function famously related to the Collatz Conjecture under the topological and ergodic perspectives, including an approach with thermodynamic formalism. By introducing a key topology and its Borel $\sigma$-algebra we show that recurrence implies periodicity. Moreover, we establish that the set of periodic orbits is finite if, and only if, every continuous potential possesses some equilibrium state. The uniqueness of periodic orbits is equivalent to the uniqueness of equilibrium state for every bounded and continuous potential. Additionally, by using the Alexandroff compactification of the established topology, we prove finiteness of cycles, which is a significant advance to the conjecture itself. We also present a proof of uniqueness of cycle and show that there is no divergent orbits. Finally, we apply our technique to the Baker and Syracuse maps, obtaining a similar result on the finiteness of orbits and no divergence for a general class of important maps.
@@ 1. Introduction
The famous Collatz function $f:\mathbb{N}\to\mathbb{N}$ is defined as follows:
@@
We intend to obtain some results on the topological and ergodic aspects of the Collatz function. It will be useful in the study of the Collatz Conjecture. We highlight its difficult by building a bridge between it and the existence and uniqueness of equilibrium states for every continuous potentials with respect to a key topology. It provides a ”dictionary” which perhaps turns possible the complete proof in another realm. We provide a significant advance by proving finiteness of cycles and later on, a proof of uniqueness of cycle. To prove the conjecture completely, we also prove that we cannot have divergent orbits. We also apply our technique for the Baker and Syracuse maps.
@@ 2. Main result
The following result is an attempt to link a very hard and famous problem on Number Theory to Ergodic Theory in the realm of Themodynamic Formalism. We list some topological and ergodic properties of the Collatz function which highlights both the beauty and the difficulty of the Collatz Conjecture using the language of equilibrium states.
@@
We have the following facts with respect to $f$
@@
  • $f$ is measurable with respect to this $\sigma$-algebra but not continuous, once $f$ is only continuous with respect to the discrete one.
@@
  • Finiteness of periodic orbits is guaranteed by the existence of an equilibrium state for every continuous potential $\phi:\mathbb{N}\to\mathbb{N}$.
@@ 3.2. Trapping the orbits towards periodicity
There exist a topology $\mathcal{T}$ and a Borel $\sigma$-algebra $\Sigma$ with respect to which the Collatz map $f:\mathbb{N}\to\mathbb{N}$ possesses $f$-invariant Borel probabilities. Moreover, every recurrent point is periodic.
@@
Let $\mathcal{M}_{f}(\mathbb{N})$ be the set of all Borelian $f$-invariant probabilities. Once there exists at least one periodic orbit, that is, $1\to 4\to 2\to 1$, we have $\mathcal{M}_{f}(\mathbb{N})\neq\emptyset$ because we can take the ergodic probability $\delta_{0}$ supported at this orbit.
Take any $\mu\in\mathcal{M}_{f}(\mathbb{N})$. By Poincaré Recurrence Theorem we have that $\mu$-almost every point $n\in\mathbb{N}$ is recurrent. Given $n_{0}\in\mathbb{N}$ a recurrent point, for every open set $n_{0}\in\mathcal{U}$ we have $f^{k}(n_{0})\in\mathcal{U}$ for some $k\in\mathbb{N}$. Hence, by taking $\mathcal{U}=\{n_{0},2n_{0}\}$ we have $f^{k}(n_{0})\in\{n_{0},2n_{0}\}$ for some $k\in\mathbb{N}$. It means that either $f^{k}(n_{0})=n_{0}$ or $f^{k}(n_{0})=2n_{0}$, that is, $f^{k+1}(n_{0})=n_{0}$. ∎
@@ 3.3. The coarseness of the key topology
3.4. Peculiarity of the discrete topology
@@ 3.5. Some key remarks
Remark 6.
The topology $\mathcal{T}$ does not contain all the singletons since it does not coincide with the discrete one by Lemma 3. The orbit $\{1,2,4\}$ is open and then a Borelian. In fact, we have the following open sets:
$\{1,2,4\}=\{1,2\}\cup\{2,4\}\quad\text{is open}.$
@@ 3.8. The ergodic decomposition
$\mu\Bigg(\bigcup_{i}\mathcal{O}_{i}\Bigg)=\sum_{i}\mu(\mathcal{O}_{i})=\sum_{i}\mu(\mathcal{O}_{i})\delta_{i}(\mathcal{O}_{i})=1.$
@@ 3.9. Finiteness of periodic orbits
Finiteness of periodic orbits is equivalent to every continuous potential $\phi:\mathbb{N}\to\mathbb{N}$ with respect to $\mathcal{T}$ possessing at least one equilibrium state.
@@
Then, there must exist finitely many periodic orbits. Otherwise, by looking at the particular potential $\varphi$ defined above, the integral $\int\varphi d\delta=\sum_{i\in\mathcal{O}(x)}i$ can attains arbitrarily large values by continuity (and measurability) and unboundedness. ∎
@@ 3.10. Uniqueness of periodic orbits: the final step
Once by hypothesis there exists a unique equilibrium state for $\chi_{\mathcal{O}}$ (which is bounded and continuous because every orbit is an open subset), we conclude that there exists a unique ergodic probability $\delta_{0}$ and the unique periodic orbit is $\{1,2,4\}$. ∎
@@
4. Finiteness of Cycles
In this section, we use the Alexandroff Compactification of topology $\mathcal{T}$ to prove finiteness of cycles. We state it as a theorem.
@@ 4. Finiteness of Cycles
Consider the Collatz map $f:\mathbb{N}\to\mathbb{N}$ and the topology $\mathcal{T}$ and the Borel $\sigma$-algebra $\Sigma$ given in Lemma 2. By using the Alexandroff Compactification of $\mathcal{T}$, we obtain finiteness of cycles for $f$.
@@
Recalling the process of compactification by Alexandroff, we know that the open sets in the compactification $\hat{\mathcal{T}}$ of the set $\hat{\mathbb{N}}$$\mathbb{N}\cup\{\infty\}$ is given by the open sets $U\in\mathcal{T}$ together with sets of the form $V=(\mathbb{N}\backslash K)\cup\{\infty\}$, where $K$ is a compact set in the topology $\mathcal{T}$. We observe that the extended map $\hat{f}$$\hat{\mathbb{N}}\to\hat{\mathbb{N}}$ is defined by
[…]
@@
Assuming, by contradiction, that there are infinitely many cycle $\mathcal{O}_{1},\cdots,\mathcal{O}_{n},\cdots$, we consider an ergodic measure $\delta_{i}$ supported at $\mathcal{O}_{i}$ and take any convex combination
$\mu:=\sum_{i}^{\infty}a_{i}\delta_{i}\quad\text{where}\quad\sum_{i}^{\infty}a_{i}=1\quad\text{and}\quad a_{i}>0.$
As we can see, the fixed point $n=\infty$ for $\hat{f}$ admits an ergodic $\delta_{\infty}$ measure supported at it, but it is distinct from the original ergodic measures $\delta_{i}$.
@@
We claim that the support of $\mu$ is exactly the union of all cycles $\{\mathcal{O}_{i}\}_{i\geq 1}$. In fact, if we had a point $x\in\text{supp}\mu$ which is not in this union, considering the open set $\{x,2x\}$ we have $\mu(\{x,2x\})>0$, which implies that $\delta_{j}(\{x,2x\})>0$ for some $j\geq 1$. It means that $\{x,2x\}\cap\mathcal{O}_{j}\neq\emptyset$, which implies that $2x\in\mathcal{O}_{j}$, that is, $f(2x)=x\in\mathcal{O}_{j}$, which is a contradiction. Then,
$\text{supp}\mu=\bigcup_{i}^{\infty}\mathcal{O}_{i}$
Once every cycle is open, it means that $\text{supp}\mu$ is open. Also, as a consequence of the definition of support, we have that $\text{supp}\mu$ is closed. So, $\text{supp}\mu$ is a clopen subset in the topology $\mathcal{T}$.
Since the open and closed subsets of $\mathcal{T}$ are also open and closed subsets in the compactification $\hat{\mathcal{T}}$, we conclude that $\text{supp}\mu$ is clopen in $\hat{\mathcal{T}}$. As every closed subset of a compact set is also compact, the subset $\text{supp}\mu$ is compact, having the union $\bigcup_{i}^{\infty}\mathcal{O}_{i}$ as a covering. By compactness, we can obtain a subcover $\bigcup_{j}^{p}\mathcal{O}_{i_{j}}$. Once the cycles are pairwise disjoint, we conclude that
$\text{supp}\mu=\bigcup_{i}^{\infty}\mathcal{O}_{i}=\bigcup_{j}^{p}\mathcal{O}_{i_{j}}.$
It means that there are at most finitely many cycles, as we wished. The theorem is proved. ∎
5. Uniqueness of Cycle
Theorem 17.
Consider the Collatz map $f:\mathbb{N}\to\mathbb{N}$ and the topology $\mathcal{T}$ and the Borel $\sigma$-algebra $\Sigma$ given in Lemma 2. The unique possible cycle is $\{1,2,4\}$.
Proof.
Let $\mathcal{O}_{1},\cdots,\mathcal{O}_{p}$ the finite list of cycles guaranteed by Theorem 16. Let us assume, by contradiction, then there exists $\mathcal{O}_{p}\neq\{1,2,4\}=\mathcal{O}_{1}$. As a consequence of the proof we know that the following set is clopen, compact and finite, with respect to the topology $\mathcal{T}$: $\mathcal{O}=\bigcup_{i=1}^{p}\mathcal{O}_{i}$.
Let $x:=\max\mathcal{O}$, which must be even because otherwise we would have $x<3x+1\in\mathcal{O}$. So $\{x,x/2\}\subset\mathcal{O}$. We have $x/2$ even because otherwise we would have $x<3(x/2)+1\in\mathcal{O}$. Also, there must exist an odd number $y\in\mathcal{O}$ such that $3y+1=x$ because otherwise $x<2x\in\mathcal{O}$. But we have the following geometric progression contained in $\mathcal{O}$: $\{y,2y,\cdots,2^{k}y\}$. But since $4y>3y+1=x=\max\mathcal{O}$, we must have $k=1$ and $\{y,2y\}\subset\mathcal{O}$. So, the following numbers are in the set $\mathcal{O}$:
$2y,y,x=3y+1,\frac{x}{2}=\frac{3y+1}{2}=y+\frac{y+1}{2}<y+y=2y.$
But we have
$y<\frac{3y+1}{2}<2y$
and the number $\frac{x}{2}=\frac{3y+1}{2}$ would be an even number in $\mathcal{O}$ between $y$ and $2y$. We also can find $z\in\mathcal{O}$ and odd number such that $3z+1=2y$. Also, the sequence $z,2z,\cdots,2^{m}z\in\mathcal{O}$. But since $8z>6z+2=2(3z+1)=2\cdot 2y=4y>x$ we must have $m=2$ and $z,2z,4z\in\mathcal{O}$. Moreover, we can find $w\in\mathcal{O}$ an odd number such that $3w+1=4z$ and $w,2w,\cdots,2^{q}w\in\mathcal{O}$. But $8w>6w+2=2(3w+1)=2\cdot 4z=8z>x$. So, we must have $q=2$ and $w,2w,4w\in\mathcal{O}$. We can find $t\in\mathcal{O}$ an odd number such that $3t+1=4w$. So $8t>6t+2=2(3t+1)=2\cdot 4w=8w>x$. By induction, we must find $a\in\mathcal{O}$ odd such that $a,2a,4a=x$. We have that
$z<y,z<w<t<\cdots<a.$
Since the cycles are pairwise disjoint, we can suppose that $\mathcal{O}_{p}$ is the unique cycle such that $x,y,z,w,t,a\in\mathcal{O}_{p}$. Moreover, the odd number $z=\min\mathcal{O}_{p}$. Then the orbit is
$\cdots\to\min\mathcal{O}_{p}=z\to 3z+1=2y\to y\to 3y+1=x=\max\mathcal{O}_{p}\to\cdots$
Then,
$x=3\frac{3z+1}{2}+1\implies x=\frac{9z+3}{2}+1\implies\max\mathcal{O}_{p}=\frac{9\min\mathcal{O}_{p}+3}{2}+1$
Also, we have $r$ geometric progressions with $3$ elements as $x=4a,2a,a$ and $1$ pair $2y,y$ and then the cycle $\mathcal{O}_{p}$ has $3r+2$ elements. We have an orbit as follows
$x=4a,2a,a,3a+1=4b,2b,b,3b+1=4c,2c,c.\cdots,4z,2z,z,3z+1=2y,y,3y+1=x.$
So, 4a4b=4a(3a+1)=a14a-4b=4a-(3a+1)=a-1, 4b4c=b1,4w4z=w14b-4c=b-1,4w-4z=w-1, which implies
$x=4a-4b+4b-4c+4c-\cdots+4w-4z+4z=$
$a-1+b-1+c-1+\cdots+w-1+z+3z=a+b+c+\cdots w+z-r+2y-1=$
$a+b+c+\cdots w+z+y-r+y-1=\frac{9z+3}{2}+1.$
The list $a,b,c,\cdots,w,z,y$ has $r+1$ odd numbers. If $r$ is odd, then $r+1$ is even and the sum $a+b+c+\cdots+w+z+y$ is even. So, yr1y-r-1 is odd and we get a contradiction because $x$ is even and the following sum is odd
$a+b+c+\cdots w+z+y-r+y-1.$
Then, $r$ must be even and we have that the sum $a+b+c+\cdots+w+z+y$ is odd. But yr1y-r-1 is even and the following sum is odd
$a+b+c+\cdots w+z+y-r+y-1.$
This contradiction shows that the cycle $\mathcal{O}_{p}$ cannot exist. We conclude that the unique cycle is $\mathcal{O}=\mathcal{O}_{1}=\{1,2,4\}$. ∎

### 18 Mar 2026 v4

@@ Abstract.
We study a class of maps having the Collatz function (famously related to the Collatz Conjecture) as an example, under the topological and ergodic perspectives, including an approach with thermodynamic formalism. By introducing a key topology and its Borel $\sigma$-algebra we show that recurrence implies periodicity. Moreover, we establish that the set of periodic orbits is finite if, and only if, every continuous potential possesses some equilibrium state. The uniqueness of periodic orbits is equivalent to the uniqueness of equilibrium state for every bounded and continuous potential. Additionally, by using the dictionary established in the paper, we prove finiteness of cycles, which is a significant advance to the conjecture itself. Finally, we apply our technique to the Baker and Syracuse maps, obtaining a similar result on the finiteness of orbits for a general class of important maps.
@@ 1. Introduction
We study a family of maps which generalizes the Collatz map, defined into two cases, namely when the natural number is either even or odd. For the case where the number $n$ is even, we always take it into its half $\frac{n}{2}$. We develop the topological and ergodic approach of this family of maps to obtain a result regarding the quantity of periodic orbits, establishing that for any of them it is finite. The quantity would be probably established by using the odd case.
Given any function $f_{0}:\mathbb{N}\to\mathbb{N}$ we can define another function $f:\mathbb{N}\to\mathbb{N}$ as follows:
[…]
As an example, we have the famous Collatz function $f:\mathbb{N}\to\mathbb{N}$ defined for $f_{0}(n)=3n+1$ as follows:
@@
We intend to obtain some results on the topological and ergodic aspects of the Collatz function. It will be useful in the study of the Collatz Conjecture. We highlight its difficult by building a bridge between it and the existence and uniqueness of equilibrium states for every continuous potentials with respect to a key topology. It provides a ”dictionary” which perhaps turns possible the complete proof in another realm. We provide a significant advance by proving finiteness of cycles. To prove the conjecture, it remains to prove that there exists a unique cycle and no divergent orbits. We also apply our technique for the Baker and Syracuse maps as follows where $a,b\in\mathbb{N}$ are odd numbers:
[…]
@@ 2. Main result
The following result is an attempt to link a very hard and famous problem on Number Theory to Ergodic Theory in the realm of Themodynamic Formalism. We list some topological and ergodic properties of the family of functions defined above which highlight both the beauty and the difficulty of the Collatz Conjecture using the language of equilibrium states.
@@
We have the following topological and ergodic facts with respect to $f$
@@
  • If $f$ the Collatz map, then it is measurable with respect to this $\sigma$-algebra but not continuous, once $f$ is only continuous with respect to the discrete one.
@@
  • Finiteness of periodic orbits is guaranteed by the existence of an equilibrium state for every continuous potential $\phi:\mathbb{N}\to\mathbb{N}$ which is integrable with respect to any $f$-invariant probability.
@@
By using Theorem A we can prove the following theorem for the general class of maps $f$, incluing the Collatz, Bake and Syracuse maps.
Theorem B.
For the family of maps defined in ()(*) there exists at most finitely many periodic orbits. Moreover, it may exists no periodic orbit.
@@ 3.2. Trapping the orbits towards periodicity
There exist a topology $\mathcal{T}$ and a Borel $\sigma$-algebra $\Sigma$ with respect to which the Collatz map $f:\mathbb{N}\to\mathbb{N}$ possesses $f$-invariant Borel probabilities. Moreover, if there exists some $f$-invariant probability, then every recurrent point is periodic.
@@
Let $\mathcal{M}_{f}(\mathbb{N})$ be the set of all Borelian $f$-invariant probabilities. Once there exists at least one $f$-invariant probability, we can take $\mu\in\mathcal{M}_{f}(\mathbb{N})$.
By Poincaré Recurrence Theorem we have that $\mu$-almost every point $n\in\mathbb{N}$ is recurrent. Given $n_{0}\in\mathbb{N}$ a recurrent point, for every open set $n_{0}\in\mathcal{U}$ we have $f^{k}(n_{0})\in\mathcal{U}$ for some $k\in\mathbb{N}$. Hence, by taking $\mathcal{U}=\{n_{0},2n_{0}\}$ we have $f^{k}(n_{0})\in\{n_{0},2n_{0}\}$ for some $k\in\mathbb{N}$. It means that either $f^{k}(n_{0})=n_{0}$ or $f^{k}(n_{0})=2n_{0}$, that is, $f^{k+1}(n_{0})=n_{0}$. ∎
Remark 3.
Once there exists some periodic orbit, we can consider some ergodic probability $\delta_{0}\in\mathcal{M}_{f}(\mathbb{N})$.
@@ 3.3. The coarseness of the key topology
3.4. Peculiarity of the discrete topology for Collatz
@@ 3.8. The ergodic decomposition
$\mu\Bigg(\bigcup_{i=1}^{\infty}\mathcal{O}_{i}\Bigg)=\sum_{i=1}^{\infty}\mu(\mathcal{O}_{i})=\sum_{i=1}^{\infty}\mu(\mathcal{O}_{i})\delta_{i}(\mathcal{O}_{i})=1.$
@@
3.9. The entropy as a convex combination
We prove that the entropy of a measure is a convex combination of the entropies of the ergodic measures, as a consequence of the ergodic decomposition.
Lemma 12.
Given
$\mu=\sum_{i=1}^{\infty}a_{i}\delta_{i},\sum_{i=1}^{\infty}a_{i}=1,$
we have
$h_{\mu}(f)=\sum_{i=1}^{\infty}a_{i}h_{\delta_{i}}(f).$
Proof.
It holds that
$\mu=\sum_{i=1}^{\infty}a_{i}\delta_{i}=\sum_{i=1}^{k}a_{i}\delta_{i}+\sum_{i=k+1}^{\infty}a_{i}\delta_{i}=\sum_{i=1}^{k}a_{i}\cdot\sum_{i=1}^{k}\frac{a_{i}}{\sum_{i=1}^{k}a_{i}}\delta_{i}+\sum_{i=k+1}^{\infty}a_{i}\cdot\sum_{i=k+1}^{\infty}\frac{a_{i}}{\sum_{i=k+1}^{\infty}a_{i}}\delta_{i}.$
Writing
$t_{k}=\sum_{i=k+1}^{\infty}a_{i},\mu_{k}=\sum_{i=1}^{k}\frac{a_{i}}{\sum_{i=1}^{k}a_{i}}\delta_{i},\nu_{k}=\sum_{i=k+1}^{\infty}\frac{a_{i}}{\sum_{i=k+1}^{\infty}a_{i}}\delta_{i},$
we obtain
$h_{\mu}(f)=h_{(1-t_{k})\mu_{k}+t_{k}\nu_{k}}(f)=(1-t_{k})h_{\mu_{k}}(f)+t_{k}h_{\nu_{k}}(f)=\sum_{i=1}^{k}a_{i}\cdot h_{\mu_{k}}(f)+t_{k}h_{\nu_{k}}(f)=$
$\sum_{i=1}^{k}a_{i}\cdot\sum_{i=1}^{k}\frac{a_{i}}{\sum_{i=1}^{k}a_{i}}h_{\delta_{i}(f)}+t_{k}h_{\nu_{k}}(f)=\sum_{i=1}^{k}a_{i}h_{\delta_{i}}(f)+t_{k}h_{\nu_{k}}(f).$
Taking $k\to\infty$, we obtain $t_{k}\to 0$ and
$h_{\mu}(f)=\sum_{i=1}^{\infty}a_{i}h_{\delta_{i}}(f).$
Remark 13.
Once we have $h_{\delta_{i}}(f)=0$ for every $i\geq 1$, then we obtain $h_{\mu}(f)=0$ for every $f$-invariant measure $\mu$.
@@ 3.10. Finiteness of periodic orbits
Finiteness of periodic orbits is equivalent to every continuous potential $\phi:\mathbb{N}\to\mathbb{N}$ with respect to $\mathcal{T}$ and integrable with respect to any $f$-invariant probability possessing at least one equilibrium state.
@@
Then, there must exist finitely many periodic orbits. Otherwise, we could have $\int\phi d\delta_{i}\to\infty$ for $\phi$ unbounded on the support. ∎
@@ 3.11. Uniqueness of periodic orbits: the final step
Once by hypothesis there exists a unique equilibrium state for $\chi_{\mathcal{O}}$ (which is bounded and continuous because every orbit is an open subset), we conclude that there exists a unique ergodic probability $\delta_{0}$ and a unique periodic orbit. ∎
@@
4. Proof of Theorem B
In this section, we use the dictionary established in Theorem A to prove the existence of finitely many cycles. We state it as a theorem.
@@ 4. Proof of Theorem B
Consider the Collatz map $f:\mathbb{N}\to\mathbb{N}$ and the topology $\mathcal{T}$ and the Borel $\sigma$-algebra $\Sigma$ given in Lemma 2. By Lemma 14, we obtain the existence of at most finitely many cycles for $f$.
@@
Assuming, by contradiction, that there are infinitely many cycles $\mathcal{O}_{1},\cdots,\mathcal{O}_{k},\cdots$, for each $i\geq 1$ we consider the ergodic measure $\delta_{i}$ supported at $\mathcal{O}_{i}$ and take any convex combination
$\mu:=\sum_{i=1}^{\infty}a_{i}\delta_{i}\quad\text{where}\quad\sum_{i=1}^{\infty}a_{i}=1\quad\text{and}\quad a_{i}>0.$
@@
We claim that the support of $\mu$ is exactly the union of all cycles $\{\mathcal{O}_{i}\}_{i\geq 1}$. In fact, if we had a point $x\in\text{supp}\mu$ which is not in this union, considering the open set $\{x,2x\}$ we have $\mu(\{x,2x\})>0$, which implies that $\delta_{j}(\{x,2x\})>0$ for some $j\geq 1$. It means that $\{x,2x\}\cap\mathcal{O}_{j}\neq\emptyset$, which implies that $2x\in\mathcal{O}_{j}$, that is, $f(2x)=x\in\mathcal{O}_{j}$, which is a contradiction. Then, $\text{supp}\mu=\bigcup_{i=1}^{\infty}\mathcal{O}_{i}$.
Given a continuous potential $\phi:\mathbb{N}\to\mathbb{N}$, integrable with respect to any $f$-invariant probability, we are going to show that there exists at least one equilibrium state.
$\sup_{\mu\in\mathcal{M}_{f}(\mathbb{N})}\left\{\int\phi d\mu\right\}=\int\phi d\delta_{i}$
In fact, we have
$\int\phi d\mu=\sum_{i=1}^{\infty}a_{i}\int\phi d\delta_{i}<\infty\implies a_{i}\int\phi d\delta_{i}\to 0.$
It is true for any convex combination and it implies that $\int\phi d\delta_{i}\leq M$ for some $M>0$. Otherwise, we could find a convex combination (integral) which diverges. In fact, if $\int\phi d\delta_{i}\to\infty$, we can take $\int\phi d\delta_{i_{j}}\geq 2^{j}$. Also, we can take $a_{i_{j}}=1/2^{j}$ and choose all the other $a_{k}$’s such that $\sum_{i=1}^{\infty}a_{i}=1$. Then,
$\int\phi d\mu=\sum_{i=1}^{\infty}a_{i}\int\phi d\delta_{i}\geq\sum_{j=1}a_{i_{j}}\int\phi d\delta_{i_{j}}\geq\sum_{j=1}^{\infty}(1/2^{j})\cdot 2^{j}=\infty.$
Once the values of the integrals are natural numbers, there must exist some $\delta_{i}$ such that it is an equilibrium state for $\phi$ by attaining the maximum integral. By Lemma 14, it means that there are finitely many cycles, as we wished. The theorem is proved. ∎